Suppose, it rains continuously at a constant rate during a thunderstorm and 1 million rain drops fall per

second, the volume of each drop is 216 mm3. If it rains continuously for 4 hours, calculate the volume of

water in litres


1
Expert's answer
2021-05-16T18:00:48-0400

"V=N\cdot t\cdot V_0=1000000\cdot(4\cdot3600)\cdot216\cdot 10^{-6}=3\ 110\ 400\ (m^2)="


"=3\ 110\ 400\ 000\ (litres)"